OpenStudy (anonymous):

In how many different ways can the top seven new indie bands be ranked on a top seven list? The top hit song for each of the seven bands will compete to receive monetary awards of $1000, $500, $250 and $100, respectively. In how many ways can the awards be given out? please show work!

OpenStudy (anonymous):

would love to know where this problem came form since this is maybe the fifth time i have seen it

OpenStudy (anonymous):

In how many different ways can the top seven new indie bands be ranked on a top seven list? answer is \[7!=7\times 6\times 5\times 4\times 3\times 2=5040\]

OpenStudy (anonymous):

lol im working on hmowrk!

OpenStudy (anonymous):

how many ways can the awards be given out? answer is \[7\times 6\times 5\times 4=840\]

OpenStudy (anonymous):

on line homework? book homework?

OpenStudy (anonymous):

book homework!